Solve the equation: x-4=5/x
Answer by Maths68(1474)

x-4=5/x
x(x-4)=5
x^2-4x=5
x^2-4x-5=0
x^2-5x+x-5=0
x(x-5)+1(x-5)=0
(x+5)(x+1)=0
x-5=0 or x+1=0
x=5 or x=-1